# The metaplectic semigroup and its applications to time-frequency analysis and evolution operators

### v1

@@ 5.1. Two subsemigroups of $\operatorname*{Mp}_{+}(d,\mathbb{C})$
where $\alpha,\beta\in\mathbb{R}$, $|\alpha|+|\beta|\neq 0$, $\beta\geq 0$ and we may assume $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$ for simplicity. The propagator of (202) consists of a one-parameter semigroup of Fourier multipliers
@@
The reader may observe that the conditions $t,\beta\geq 0$ not only guarantee the well-posedness of (202) for $t\geq 0$, but also the validity of the Hörmander condition of Definition 3.1 for $V_{-P_{t}}^{\top}$. Stated differently, $V_{-P_{t}}^{\top}\in\operatorname*{Sp}_{+}(d,\mathbb{C})$ and $\mathfrak{m}_{P_{t}}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ for every $t\geq 0$.
@@
Recall the definition of $\mathcal{M}_{S}$ in (104). Let us write $A=A_{1}+iA_{2}$, $C=C_{1}+iC_{2}$ and $A^{-\top}=D_{1}+iD_{2}$. Of course, $A\in\operatorname*{GL}(d,\mathbb{R})$ if and only if $A_{2}=O$ or, equivalently, $D_{2}=O$. A straightforward computation allows us to write
(208) $\mathcal{M}_{S}=\mathcal{K}\left(\begin{array}[]{cc|cc}-C_{1}^{\top}A_{2}+A_{1}^{\top}C_{2}&-C_{2}^{\top}A_{2}+A_{2}^{\top}C_{2}&A_{1}^{\top}D_{2}&A_{2}^{\top}D_{2}\\ C_{2}^{\top}A_{2}-A_{2}^{\top}C_{2}&-C_{1}^{\top}A_{2}+A_{1}^{\top}C_{2}&-A_{2}^{\top}D_{2}&A_{1}^{\top}D_{2}\\ \hline\cr-D_{1}^{\top}A_{2}&-D_{2}^{\top}A_{2}&O&O\\ D_{2}^{\top}A_{2}&-D_{1}^{\top}A_{2}&O&O\end{array}\right)\mathcal{K},$
@@
(211) $\begin{pmatrix}A_{1}^{\top}C_{2}&O\\ O&A_{1}^{\top}C_{2}\end{pmatrix}\geq 0$
@@ 5.2. Wigner analysis
In this subsection, we compute the Wigner operator of a general $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ and, by the above argument, this amounts to computing $K^{(2)}$ in (215).
As in the previous sections, if $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$, we use the factorization (124) and write $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$, with $\widehat{U}_{1,2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ atomic metaplectic contraction. For the rest of this section, we omit the phase factors defining metaplectic operators. We also recall that
@@
Under the notation above,
@@
We need to compute $K_{\Xi}$. The following lemma is proven in [12].
@@
(226) $\mathcal{A}_{1/2}=\begin{pmatrix}I_{d}/2&I_{d}/2&O_{d}&O_{d}\\ O_{d}&O_{d}&I_{d}/2&-I_{d}/2\\ O_{d}&O_{d}&I_{d}&I_{d}\\ -I_{d}&I_{d}&O_{d}&O_{d}\end{pmatrix}$
@@
The equivalent of Lemma 5.5 was proven in [12].
@@
Equations (223) and (232) together give the operator $K^{\prime}$ in (221).
@@
(245) $=(\mathcal{F}_{2}\mathfrak{T}_{2^{-1/2}I_{2d}})^{-1}(\widehat{S}_{2}\otimes\widehat{S}_{2})(\mathcal{F}_{2}\mathfrak{T}_{2^{-1/2}I_{2d}})W(f,g),$
@@ 5.3. Metaplectic operators and complex conjugation
In this subsection, we delve into the details of the intertwining relation between metaplectic operators and complex conjugation, thereby rephrasing [13, Appendix A] in our context and proving Theorem 1.7. The characterization of metaplectic operators which commute with complex conjugation is then provided in terms of the structure of their projections onto $\operatorname*{Sp}_{+}(d,\mathbb{C})$.
@@
As always, let us write $S=U_{1}\Xi U_{2}$ with $U_{1,2}\in\operatorname*{Sp}(d,\mathbb{R})$ and $\Xi$ metaplectic atomic contraction. Since Gaussian products trivially commute with complex conjugation and also the atoms $\mathfrak{R}_{\Theta}$ commute with complex conjugation, as it can be verified directly from (116), we have by (132) that $\widehat{\Xi}\bar{f}=\overline{\widehat{\Xi}f}$. By applying Proposition 26, we get
@@
Since $\widehat{\widetilde{U}}_{1,2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ is an atomic metaplectic contraction, $T=\widehat{\widetilde{U}}_{1}\widehat{\Xi}\widehat{\widetilde{U}}_{2}\in\operatorname*{Mp}_{+}(d,\mathbb{R})$. Its projection is $\widetilde{S}=\widetilde{U}_{1}\Xi\widetilde{U}_{2}$. By writing explicitly $\Xi$ as in (132), we see that $\Re(B_{2})=\Re(C_{2})=O_{d}$ and $\Im(A_{2})=\Im(D_{2})=O_{d}$, therefore if
(249) $S_{j}=\begin{pmatrix}A_{j}&B_{j}\\ C_{j}&D_{j}\end{pmatrix},\qquad j=1,2,3,$
@@
    $R(C^{\top}B)\subseteq R(C^{\top}A)$.
@@
$(ii^{\prime})$ $R(B)\subseteq R(A)$.
The implication $(ii^{\prime})\Rightarrow(i)$ is trivial, and thus we prove the converse. Assume that $R(C^{\top}B)\subseteq R(C^{\top}A)$. For $\zeta=Bx\in R(B)$, $x\in\mathbb{R}^{d}$, we must show that there exists $z\in\mathbb{R}^{d}$ such that $\zeta=Az$. By the assumption, there exists $y\in\mathbb{R}^{d}$ such that $C^{\top}Bx=C^{\top}Ay$. Then, there exists $v\in\ker(C^{\top})=R(C)^{\perp}$ such that BxAy=vBx-Ay=v. By Proposition 2.4 $(i)$, there exists a unique $w\in\ker(C)$ such that $Aw=v$. It follows that $Bx=A(y+w)\in R(A)$. We now turn to the equivalence $(ii^{\prime})\Leftrightarrow(iii)$. Observe that if $A$ is invertible, then $(ii^{\prime})$ is trivial. We need to check that $(ii^{\prime})\Rightarrow(iii)$. If $R(B)\subseteq R(A)$, then $R(A)^{\perp}\subseteq R(B)^{\perp}=\ker(B^{\top})$. By item $(ii)$ of Proposition 2.4, we obtain $\ker(A)=B^{\top}(R(A)^{\perp})=\{0\}$. This concludes the proof.
We are ready to prove the main result of this section.
@@
where $\mathcal{K}$ is defined as in (209). By Theorem 2.1, condition (262) is equivalent to simultaneous validity of the following:
@@
    $R(C^{\top}B)\subseteq R(C^{\top}A)$.
@@
Condition $R(C^{\top}B)\subseteq R(C^{\top}A)$ is equivalent either to $R(B)\subseteq R(A)$ or to $A\in\operatorname*{GL}(d,\mathbb{R})$ by Lemma 5.8 applied to $S^{\prime}$ in (257). On the other hand, item (3) is equivalent to
@@ 6. Applications to time-frequency analysis
The notion of time-frequency representation lies at the foundation of time-frequency analysis. Across essentially every line of inquiry in the field, one is naturally led to metaplectic operators. However, it turns out that the structure of $\operatorname*{Mp}(d,\mathbb{R})$ is not rich enough to include regularizing operators, such as Gaussian filtering and the propagators of heat equations. In the present section, we define the analogue of metaplectic Wigner distributions in the context of the metaplectic semigroup $\operatorname*{Mp}_{+}(2d,\mathbb{C})$, and we characterize those that satisfy certain fundamental properties arising in operator theory, time-frequency analysis and quantum mechanics.
@@
The reader may observe that this definition resembles [18, Definition 4.1], with the difference that we now allow $\widehat{\mathcal{A}}\in\operatorname*{Mp}_{+}(2d,\mathbb{C})$. The phraseologies metaplectic Wigner distribution and $\mathcal{A}$-Wigner distribution are used in the aforementioned work, and in the subsequent ones by the same authors, in the case $\widehat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{R})$. This shall not cause confusion; therefore, we adopt the same terminology in our more general setting to avoid unnecessary burden on notation and terminology.
@@
We need the following lemma.
@@
The first part of the Lemma is trivial, we prove it for the sake of completeness. If $\Im(Q)\geq 0$, the Gaussian $e^{-\pi\Im(Q)x\cdot x}=1$ if and only if $x\in\ker(Q)$. Let $\Gamma=\{y:\mathrm{dist}(x,\ker(Q))<1\}$. Then,
@@
(280) $\mu_{j}=\begin{cases}\alpha_{j}&\text{if $\widehat{S}_{j}=\mathfrak{p}_{i\alpha_{j}}$},\\ \vartheta_{j}&\text{if $\widehat{S}_{j}=\mathfrak{R}_{\vartheta_{j}}$},\\ 0&\text{otherwise}.\end{cases}$
@@
As detailed in the following sections, the transition from metaplectic Wigner distributions $W_{\mathcal{U}}$ with $\widehat{\mathcal{U}}\in\mathrm{Mp}(2d,\mathbb{R})$ to those $W_{\mathcal{A}}$ with $\widehat{\mathcal{A}}\in\mathrm{Mp}_{+}(2d,\mathbb{C})$ provides distributions that satisfy several properties which, in the real case, would only hold for a very restricted subclass of metaplectic Wigner distributions. Continuity properties of metaplectic Wigner distributions follow by those of the operators in $\operatorname*{Mp}_{+}(d,\mathbb{C})$. However, this comes at the cost of losing the isometry property on $L^{2}$ and Moyal’s identity, as the next two results clarify.
@@
Let $W_{\mathcal{A}}$ be a a complex metaplectic Wigner distribution.
@@
Items $(i)$ and $(iii)$, along with the $L^{2}$ boundedness in $(ii)$, follow simply by the continuity properties of operators in $\operatorname*{Mp}_{+}(2d,\mathbb{C})$. The only item that remains to verify is the “only if” part of $(ii)$. To prove it, we check that if $W_{\mathcal{A}}$ has $\widehat{\mathcal{A}}\in\operatorname*{Mp}_{+}(d,\mathbb{C})\setminus\operatorname*{Mp}(d,\mathbb{R})$, then there exist functions $f,g\in L^{2}(\mathbb{R}^{d})$ so that (282) holds with a strict inequality. By Lemma 4.4, we may decompose $\widehat{\mathcal{A}}=\widehat{\mathcal{U}}\widehat{\mathcal{Z}}$ with $\widehat{\mathcal{U}}\in\operatorname*{Mp}(2d,\mathbb{R})$ and $\widehat{\mathcal{Z}}\in\operatorname*{Mp}_{0}(2d,\mathbb{C})$, and $\widehat{\mathcal{Z}}$ is not a multiple of the identity. The left-hand side of (282) reads as
@@
The “if” part is trivial. The “only if” part follows then by Proposition 6.4, since Moyal’s identity implies the equality in (282), holding if and only if $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})$.

### v2

@@ 5.1. Two subsemigroups of $\operatorname*{Mp}_{+}(d,\mathbb{C})$
where $\alpha,\beta\in\mathbb{R}$, $|\alpha|+|\beta|\neq 0$, $\beta\geq 0$ and $u_{0}\in\mathcal{S}(\mathbb{R}^{d})$. The propagator of (204) consists of a one-parameter semigroup of Fourier multipliers
@@
The reader may observe that the conditions $t,\beta\geq 0$ not only guarantee the well-posedness of (204) for $t\geq 0$, but also the validity of the condition of Definition 3.1 for $V_{-P_{t}}^{\top}$. Stated differently, $V_{-P_{t}}^{\top}\in\operatorname*{Sp}_{+}(d,\mathbb{C})$ and $\mathfrak{m}_{P_{t}}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ for every $t\geq 0$.
@@
Recall the definition of $\mathcal{M}_{S}$ in (103). Let us write $A=A_{1}+iA_{2}$, $C=C_{1}+iC_{2}$ and $A^{-\top}=D_{1}+iD_{2}$. Of course, $A\in\operatorname*{GL}(d,\mathbb{R})$ if and only if $A_{2}=O_{d}$ or, equivalently, $D_{2}=O_{d}$. A straightforward computation allows us to write
(210) $\mathcal{M}_{S}=\mathcal{K}\left(\begin{array}[]{cc|cc}-C_{1}^{\top}A_{2}+A_{1}^{\top}C_{2}&-C_{2}^{\top}A_{2}+A_{2}^{\top}C_{2}&A_{1}^{\top}D_{2}&A_{2}^{\top}D_{2}\\ C_{2}^{\top}A_{2}-A_{2}^{\top}C_{2}&-C_{1}^{\top}A_{2}+A_{1}^{\top}C_{2}&-A_{2}^{\top}D_{2}&A_{1}^{\top}D_{2}\\ \hline\cr-D_{1}^{\top}A_{2}&-D_{2}^{\top}A_{2}&O_{d}&O_{d}\\ D_{2}^{\top}A_{2}&-D_{1}^{\top}A_{2}&O_{d}&O_{d}\end{array}\right)\mathcal{K},$
@@
(213) $\begin{pmatrix}A_{1}^{\top}C_{2}&O_{d}\\ O_{d}&A_{1}^{\top}C_{2}\end{pmatrix}\geq 0$
@@ 5.2. Wigner analysis
In this subsection, we compute the Wigner operator of a general $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$ and, by the above argument, this amounts to computing $K_{Z}$ in (217).
As in the previous sections, if $\widehat{S}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$, we use the factorization (123) and write $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$, with $\widehat{U}_{1},\widehat{U}_{2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ atomic metaplectic contraction. For the rest of this section, we omit the phase factors defining metaplectic operators. We also recall that
@@
Under the notation above, $\widehat{S}=\widehat{U}_{1}\widehat{\Xi}\widehat{U}_{2}$,
@@
We need to compute $K_{\Xi}$. The following lemma is proven in a recent, unpublished, manuscript by two of the authors  [^2] See E. Cordero, G. Giacchi, and L. Rodino. Wigner and Gabor phase-space analysis of propagators for evolution equations. arXiv:2511.19400..
@@
(228) $\mathcal{A}_{1/2}=\begin{pmatrix}I_{d}/2&I_{d}/2&O_{d}&O_{d}\\ O_{d}&O_{d}&I_{d}/2&-I_{d}/2\\ O_{d}&O_{d}&I_{d}&I_{d}\\ -I_{d}&I_{d}&O_{d}&O_{d}\end{pmatrix}$.
@@
The equivalent of Lemma 5.6 was proven in the unpublished work in Footnote 2.
@@
Equations (225) and (234) together give the operator $K_{\Xi}$ in (223).
@@
(247) $=(\mathcal{F}_{2}\mathfrak{T}_{2^{-1/2}I_{2d}})^{-1}(\widehat{\Xi}\otimes\widehat{\Xi})(\mathcal{F}_{2}\mathfrak{T}_{2^{-1/2}I_{2d}})W(f,g),$
@@ 5.3. Metaplectic operators and complex conjugation
In this subsection, we delve into the details of the intertwining relation between metaplectic operators and complex conjugation, thereby rephrasing [12, Appendix A] in our context and proving Theorem 1.7. The characterization of metaplectic operators which commute with complex conjugation is then provided in terms of the structure of their projections onto $\operatorname*{Sp}_{+}(d,\mathbb{C})$.
@@
As always, let us write $S=U_{1}\Xi U_{2}$ with $U_{1},U_{2}\in\operatorname*{Sp}(d,\mathbb{R})$ and $\Xi$ metaplectic atomic contraction. Since Gaussian products trivially commute with complex conjugation and also the atoms $\mathfrak{R}_{\Theta}$ commute with complex conjugation, as it can be verified directly from (115), we have by (131) that $\widehat{\Xi}\bar{f}=\overline{\widehat{\Xi}f}$. By applying Proposition 1.6, we get
@@
Since $\widehat{\widetilde{U}}_{1,2}\in\operatorname*{Mp}(d,\mathbb{R})$ and $\widehat{\Xi}$ is an atomic metaplectic contraction, $T=\widehat{\widetilde{U}}_{1}\widehat{\Xi}\widehat{\widetilde{U}}_{2}\in\operatorname*{Mp}_{+}(d,\mathbb{C})$. Its projection is $\widetilde{S}=\widetilde{U}_{1}\Xi\widetilde{U}_{2}$. By writing explicitly $\Xi$ as in (131), we see that $\Re(B_{2})=\Re(C_{2})=O_{d}$ and $\Im(A_{2})=\Im(D_{2})=O_{d}$, therefore if
(251) $U_{1}=\begin{pmatrix}A_{1}&B_{1}\\ C_{1}&D_{1}\end{pmatrix},\qquad\Xi=\begin{pmatrix}A_{2}&B_{2}\\ C_{2}&D_{2}\end{pmatrix},\qquad U_{2}=\begin{pmatrix}A_{3}&B_{3}\\ C_{3}&D_{3}\end{pmatrix}$,
@@
    $\mathrm{range}(C^{\top}B)\subseteq\mathrm{range}(C^{\top}A)$.
@@
$(ii^{\prime})$ $\mathrm{range}(B)\subseteq\mathrm{range}(A)$.
The implication $(ii^{\prime})\Rightarrow(i)$ is trivial, and thus we prove the converse. Assume that $\mathrm{range}(C^{\top}B)\subseteq\mathrm{range}(C^{\top}A)$. For $\zeta=Bx\in\mathrm{range}(B)$, $x\in\mathbb{R}^{d}$, we must show that there exists $z\in\mathbb{R}^{d}$ such that $\zeta=Az$. By the assumption, there exists $y\in\mathbb{R}^{d}$ such that $C^{\top}Bx=C^{\top}Ay$. Then, there exists $v\in\ker(C^{\top})=\mathrm{range}(C)^{\perp}$ such that BxAy=vBx-Ay=v. By Proposition 2.4 $(i)$, there exists a unique $w\in\ker(C)$ such that $Aw=v$. It follows that $Bx=A(y+w)\in\mathrm{range}(A)$. We now turn to the equivalence $(ii^{\prime})\Leftrightarrow(iii)$. Observe that if $A$ is invertible, then $(ii^{\prime})$ is trivial. We need to check that $(ii^{\prime})\Rightarrow(iii)$. If $\mathrm{range}(B)\subseteq\mathrm{range}(A)$, then $\mathrm{range}(A)^{\perp}\subseteq\mathrm{range}(B)^{\perp}=\ker(B^{\top})$. By item $(ii)$ of Proposition 2.4, we obtain $\ker(A)=B^{\top}(\mathrm{range}(A)^{\perp})=\{0\}$. This concludes the proof.
We are ready to prove the main result of this subsection.
@@
where $\mathcal{K}$ is defined as in (211). By Theorem 2.1, condition (264) is equivalent to the simultaneous validity of the following:
@@
    $\mathrm{range}(C^{\top}B)\subseteq\mathrm{range}(C^{\top}A)$.
@@
Condition $\mathrm{range}(C^{\top}B)\subseteq\mathrm{range}(C^{\top}A)$ is equivalent either to $\mathrm{range}(B)\subseteq\mathrm{range}(A)$ or to $A\in\operatorname*{GL}(d,\mathbb{R})$ by Lemma 5.9 applied to $S^{\prime}$ in (259). On the other hand, item (3) is equivalent to
@@ 6. Applications to time-frequency analysis
The notion of time-frequency representation lies at the foundation of time-frequency analysis. Across essentially every line of inquiry in the field, one is naturally led to metaplectic operators. In the present section, we define the analogue of metaplectic Wigner distributions in the context of the metaplectic semigroup $\operatorname*{Mp}_{+}(2d,\mathbb{C})$, and we characterize those that satisfy certain fundamental properties arising in operator theory, time-frequency analysis and quantum mechanics.
@@
The reader may observe that this definition resembles [17, Definition 4.1], with the difference that we now allow $\widehat{\mathcal{A}}\in\operatorname*{Mp}_{+}(2d,\mathbb{C})$. The phraseologies metaplectic Wigner distribution and $\mathcal{A}$-Wigner distribution are used in the aforementioned work, and in the subsequent ones by the same authors, in the case $\widehat{\mathcal{A}}\in\operatorname*{Mp}(2d,\mathbb{R})$. This shall not cause confusion; therefore, we adopt the same terminology in our more general setting to avoid unnecessary burden on notation and terminology.
@@
We need the following lemma, whose proof is reported for the sake of completeness.
@@
If $\Im(Q)\geq 0$, the Gaussian $e^{-\pi\Im(Q)x\cdot x}=1$ if and only if $x\in\ker(Q)$. Let $\Gamma=\{y:\mathrm{dist}(y,\ker(Q))<1\}$. Then,
@@
(282) $\mu_{j}=\begin{cases}\alpha_{j}&\text{if $\widehat{\Xi}_{j}=\mathfrak{p}_{i\alpha_{j}}$},\\ \vartheta_{j}&\text{if $\widehat{\Xi}_{j}=\mathfrak{R}_{\vartheta_{j}}$},\\ 0&\text{otherwise}.\end{cases}$
@@
As detailed in the following sections, the transition from metaplectic Wigner distributions $W_{\mathcal{U}}$ with $\widehat{\mathcal{U}}\in\mathrm{Mp}(2d,\mathbb{R})$ to those $W_{\mathcal{A}}$ with $\widehat{\mathcal{A}}\in\mathrm{Mp}_{+}(2d,\mathbb{C})$ provides distributions that satisfy several properties which, in the real case, would only hold for a very restricted, and basically trivial, subclass of metaplectic Wigner distributions. Continuity properties of metaplectic Wigner distributions follow by those of the operators in $\operatorname*{Mp}_{+}(d,\mathbb{C})$. However, this comes at the cost of losing the isometry property on $L^{2}$ and Moyal’s identity, as the next two results clarify.
@@
Let $W_{\mathcal{A}}$ be a a metaplectic Wigner distribution with $\mathcal{A}\in\operatorname*{Sp}_{+}(2d,\mathbb{C})$.
@@
Items $(i)$ and $(iii)$, along with the $L^{2}$ boundedness in $(ii)$, follow simply by the continuity properties of operators in $\operatorname*{Mp}_{+}(2d,\mathbb{C})$. The only item that remains to verify is the “only if” part of $(ii)$. To prove it, we check that if $W_{\mathcal{A}}$ has $\widehat{\mathcal{A}}\in\operatorname*{Mp}_{+}(d,\mathbb{C})\setminus\operatorname*{Mp}(d,\mathbb{R})$, then there exist functions $f,g\in L^{2}(\mathbb{R}^{d})$ so that (284) holds with a strict inequality. By Theorem 4.4, we may decompose $\widehat{\mathcal{A}}=\widehat{\mathcal{U}}\widehat{\mathcal{Z}}$ with $\widehat{\mathcal{U}}\in\operatorname*{Mp}(2d,\mathbb{R})$ and $\widehat{\mathcal{Z}}\in\operatorname*{Mp}_{0}(2d,\mathbb{C})$, and $\widehat{\mathcal{Z}}$ is not a multiple of the identity. The left-hand side of (284) reads as
@@
The “if” part is trivial. The “only if” part follows then by Proposition 6.4, since Moyal’s identity implies the equality in (284), which holds if and only if $\mathcal{A}\in\operatorname*{Sp}(2d,\mathbb{R})$.